# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Problem-Space Evasion Adversarial Attacks on Flow-based ML-NIDS

### v3

@@ Abstract
As network security threats continue to evolve, safeguarding flow-based Machine Learning (ML)-based Network Intrusion Detection Systems (NIDS) from adversarial attacks is crucial. This paper introduces the notion of feature perturb-ability and presents a novel Perturb-ability Score (PS) metric, which identifies NIDS features that are susceptible to manipulation in the problem-space by an attacker. By quantifying a feature’s susceptibility to perturbations within the problem-space, the PS facilitates the selection of features that are inherently more robust against evasion adversarial attacks on flow-based ML-NIDS during the feature selection phase. These features exhibit natural resilience to perturbations as they are heavily constrained by the problem-space limitations and correlations specific to the NIDS domain. Furthermore, manipulating these features may either disrupt the malicious functionality of evasion adversarial attacks on NIDS or render the network traffic invalid for processing (or both). This novel approach introduces a fresh perspective by leveraging problem-space NIDS domain constraints as a defense mechanism against problem-space evasion adversarial attacks targeting flow-based ML-NIDS.
We demonstrate the effectiveness of our PS-guided feature selection defense in enhancing flow-based NIDS robustness. Experimental results across various ML-based NIDS models and public datasets show that selecting only robust features (low-PS features) can maintain solid detection performance while significantly reducing vulnerability to evasion adversarial attacks. Additionally, our findings verify that the PS effectively identifies flow-based NIDS features highly vulnerable to problem-space perturbations.
[…]
L1/.style=fill=white,, L2/.style=fill=myred,edge=myred,line width=2pt, L3/.style=fill=yellow,edge=yellow,line width=2pt, L4/.style=fill=mygreen,edge=mygreen,line width=2pt,
1 Introduction
@@
(b) Two Examples of Non-Perturb-able Features
in Problem-Space Against NIDS
@@
1.1 Feature-Space vs Problem-Space Evasion Adversarial Attacks Against ML-NIDS
It is crucial to differentiate between feature-space and problem-space (real-world objects) in ML [10]. Feature-space refers to the representation of input data as features used by the ML model, while problem-space refers to the raw input data before feature extraction. Ibitoye et al. [25] introduced the concept of ’space’ in their taxonomy of adversarial attacks for network security, distinguishing between feature-space and problem-space attacks. Feature-space adversarial attacks manipulate or perturb feature vectors, assuming that the attacker can directly access and alter these features; however, this assumption is often unrealistic in practical scenarios. In contrast, problem-space adversarial attacks modify or perturb actual raw input data, such as network traffic flows (e.g., adding delays) or packets (e.g., padding the payload), which attackers can typically access. Fig. 1 illustrates this distinction.
Feature-space attacks may not be practical against NIDS due to challenges an attacker would face in feature vector access and challenges with feature correlations and network constraints [35]. On the other hand, problem-space attacks are more practical than feature-space as the modifications happen to the network packets (feasible for an external attacker). They typically start with feature-space perturbations, then translate to real-world packet modifications (Inverse Feature-Mapping [31]). In other words, as seen in Fig. 2, attackers modify the adversarial network flow to produce perturbations in certain targeted features within the feature vector after feature extraction. Fig. 2 illustrates an example of evasion adversarial attack problem-space perturbations against NIDS, where the attacker adds padding to the payload of a packet to perturb features such as the maximum packet length and total length of forward packets in the feature vector. Despite being considered more practical than feature-space attacks, these attacks also face several practicality issues [35] [20] [6], such as: challenges in maintaining malicious objective and network functionality while altering packets; keeping up-to-date knowledge of the model, its features, and extraction techniques; or predicting correct common features. Problem-space attacks must also address NIDS features constraints.
1.2 Perturb-ability of Features in Problem-Space Against NIDS
@@
N.B. Some non-perturb-able (or robust) features may be completely unmodifiable due to problem-space limitations and correlations within the NIDS domain. For example, some backward features (features describing the network flow from server to client), like the mean size of a packet in the backward direction feature, can be extremely difficult for an attacker to access. However, it is important to note that most features can be modified through problem-space manipulations. By non-perturb-able features, we specifically refer to those that cannot be perturbed in the problem space while maintaining the attacker’s malicious aim and complying with NIDS domain constraints. For instance, changing the destination IP to manipulate its corresponding feature in the feature vector is possible. However, doing so would disrupt the flow’s network functionality and the attacker’s malicious objective, which is why we classify it as a non-perturb-able feature.
@@ 1.3 Motivation and Aim
Our motivation stems from the intuitive assumption that attackers can only access the problem-space rather than the feature-space. This perspective aligns with the reality of most network environments, where attackers can manipulate packet contents but do not have direct control over the feature extraction process (see section 2 for more details on our threat model).
@@
In response to this, our aim is to introduce the novel notion of the Perturb-ability Score (PS) metric, which is designed to enhance the robustness of ML-based NIDS. The PS metric helps to identify features in the problem-space that are susceptible to manipulation by attackers, without compromising the malicious functionality of network traffic. By quantifying the perturb-ability of each feature within NIDS domain constraints, PS facilitates the selection of features that are inherently more resistant to adversarial attacks. Our aimed classification is shown in Fig. 4. For the remainder of the paper, we will use the color scheme found in Fig. 4, i.e., a green feature represents a feature in the Low Perturb-ability class, a yellow feature represents a feature in the Medium Perturb-ability class, and a red feature represents a feature in the High Perturb-ability class.
Our approach embodies this philosophy by offering a straightforward yet highly effective defense against evasion adversarial attacks targeting ML-based Network Intrusion Detection Systems (ML-NIDS). What sets our solution apart is its independence from attack types, attack-norms [11] used, or the level of adversarial knowledge (whether black-box, white-box, or gray-box [25]).
@@
Unlike most conventional defenses that focus primarily on the internal mechanisms of the ML model itself our method takes an “outside-the-box” perspective. By leveraging inherent network domain constraints external to the ML model, we significantly reduce the attack surface. This strategic shift from model-centric defenses to exploiting network domain properties introduces a novel layer of protection, enhancing the overall robustness of ML-NIDS against adversarial threats.
In this paper, we demonstrate the efficacy of PS-guided feature selection in fortifying ML-based NIDS against adversarial attacks. Through extensive experiments on various NIDS models and public datasets, we show that PS-selected features maintain solid detection performance while significantly reducing vulnerability to adversarial manipulations. By “maintain solid detection performance,” we mean that the difference in performance metrics between the ML-NIDS model trained exclusively on non-perturb-able (robust) features and the ML-NIDS model trained on all features is minor.
This work represents a promising step towards developing resilient NIDS capable of withstanding the evolving threat landscape.
1.4 Contributions
@@
Our contributions are threefold: (1) Introduction of Perturb-ability Score (PS) (section 3): We propose a novel the Perturb-ability Score (PS) metric to quantify the vulnerability of NIDS features to adversarial manipulation. The PS measures the susceptibility of a feature to perturbation by an attacker without compromising the underlying malicious objective or network functionality of the attack. This metric establishes a robust foundation for assessing feature robustness in ML-NIDS adversarial contexts. (2) Leveraging PS as a Defensive Mechanism (section 3.4): By utilizing PS, we introduce a defense strategy that enhances the robustness of ML-based NIDS. PS-guided feature selection enables the identification of inherently resilient features that remain robust against adversarial perturbations. This approach fortifies NIDS by reducing the attack surface while maintaining solid detection performance across various models and datasets. (3) Mapping Problem-Space Adversarial Attacks to feature-space (section 5.4): We conduct an in-depth analysis of problem-space adversarial attack techniques found in literature by mapping their traffic morphing methods to the corresponding NIDS features. This mapping validates the PS classification, demonstrating how PS effectively captures the impact of problem-space evasion techniques on NIDS features and provides significant insights into how adversarial attacks manifest in both problem and feature-spaces.
2 Threat Model
@@
Although recent research has cast doubt on the practicality of evasion adversarial attacks against NIDS [35, 7, 16], since knowing the ML algorithm, activation functions, or querying the NIDS is unlikely, this paper nonetheless focuses on exploring the possibility and implications of adding perturbations to NIDS features in the problem space.
3 Perturb-ability Score (PS)
In this section, we explain how our novel Perturbability Score (PS) quantifies the susceptibility of each feature to perturbation and facilitates defense against practical problem-space evasion attacks.
3.1 NIDS Features Perturb-ability Classification
@@ 3.2 Evaluating PS
(a) Visualization of the PS2[$f_{i}$] Evaluation Equation
@@
3.2.1 PS1[$f_{i}$]: Strict Header Features/ Affects Network or Malicious Functionality
This PS field focuses on strict Header features and network/malicious functionality of network flows after adding perturbations in the problem-space. PS1[$f_{i}$] will be 0 if any of the following conditions are true (which will make PSTotal[$f_{i}$] equals 0);
C1: the feature $f_{i}$ is a strict header feature (IP addresses in TCP flows, destination port number or protocol)
C2: adding perturbation to feature $f_{i}$ will affect the network or malicious functionality of the flow.
@@ 3.2.1 PS1[$f_{i}$]: Strict Header Features/ Affects Network or Malicious Functionality
3.2.2 PS2[$f_{i}$]: The range of Possible Values of a Feature
@@
PS2[$f_{i}$] will be 1 if $f_{i}$’s number of Possible Values ($PV$) is greater than 255 (this feature will be similar to computer vision’s pixel, and it will be flexible to perturb). On the other hand, if $f_{i}$’s $PV$ ($PV$[$f_{i}$]) is less than or equal to 255, PS2[$f_{i}$] will be equal to a linear function where its output is 1 if $f_{i}$’s $PV$ is 255, and 0.5 if $f_{i}$’s $PV$ is 2 (binary). If $PV[f_{i}]$ is less than 2 (equals 1), it indicates that $f_{i}$ is non-perturb-able, in which case $\text{PS}_{2}[f_{i}]$ will be set to 0. However, in this case (where $PV[f_{i}]$ equals 1), we recommend dropping that feature, as it does not contribute meaningful information to the ML model. Fig. 5a shows the visualization of the PS2[$f_{i}$] evaluation equation.
@@
$\text{PS}_{2}[f_{i}]=\begin{cases}1&\text{if }\text{PV[$f_{i}$]}>255\\ 0&\text{if }\text{PV[$f_{i}$]}<2\\ 0.5+(0.5\times\frac{(\text{PV[$f_{i}$]}-2)}{(255-2)})&\text{otherwise}\end{cases}$
@@
3.2.3 PS3[$f_{i}$]: Correlated Features
This PS field considers the correlation between a NIDS feature and other features. Due to network constraints within NIDS, many features exhibit problem-space correlations. For instance, the flow duration feature is typically correlated with the total forward and backward inter-arrival times. Such correlated features limit the attacker’s flexibility. The gradients of the targeted model might recommend a specific perturbation to one feature and a different perturbation to another. However, achieving these opposing perturbations simultaneously is very difficult if the features are highly correlated within the problem-space. As an example, an attacker cannot simultaneously increase the flow duration while decreasing both the forward and backward inter-arrival times. As the number of correlated features associated with a single feature increases, it becomes more difficult to perturb that feature in the problem-space.
PS3[$f_{i}$] will follow a linear function, where its output is 0.5 if the number of Correlated Features ($CF$) of $f_{i}$ is equal to or greater than a threshold (the maximum number observed in our experiments was 10, which we chose as the threshold), and 1 if $f_{i}$’s $CF$ (CF[$f_{i}$]) is 0. Fig. 5b shows the visualization of the PS3[$f_{i}$] evaluation equation.
PS3[$f_{i}$] can be described with the following equation:
$\text{PS}_{3}[f_{i}]=1-0.05\times\min(\text{CF[$f_{i}$]},10)$
@@ 3.2.3 PS3[$f_{i}$]: Correlated Features
Nevertheless, we acknowledge that an attacker could potentially overcome the constraints posed by correlated features in the problem-space, provided they have exceptional knowledge of the attacked ML model and its feature vector. To address this, we implemented a linear function that slightly reduces PS3[$f_{i}$] in the case of two correlated features associated with a single feature. As the number of correlations increases, the function output decreases until it reaches 0.5 (not 0) in the case of 10 correlations, which is the maximum number we have observed. In the worst-case scenario for an attacker (10 correlations or more per feature), the value of PS3[$f_{i}$] will only drop to 0.5, resulting in a modest reduction in PStotal[$f_{i}$]. However, we still believe that correlated features in the problem-space would pose a significant challenge for an attacker in a practical setting where they have no information about the attacked ML-NIDS and cannot query it. Yet, we opted to only decrease the PStotal[$f_{i}$] value, rather than nullify it, to account for the possibility of more advanced future attacks in the problem-space.
3.2.4 PS4[$f_{i}$]: Features that attackers cannot access

### v4

# A Novel Perturb-ability Score to Mitigate Evasion Adversarial Attacks on Flow-Based ML-NIDS

@@ Abstract
As network security threats evolve, safeguarding flow-based Machine Learning (ML)-based Network Intrusion Detection Systems (NIDS) from evasion adversarial attacks is crucial. This paper introduces the notion of feature perturb-ability and presents a novel Perturb-ability Score (PS), which quantifies how susceptible NIDS features are to manipulation in the problem-space by an attacker. PS thereby identifies features structurally resistant to evasion attacks in flow-based ML-NIDS due to the semantics of network traffic fields, as these features are constrained by domain-specific limitations and correlations. Consequently, attempts to manipulate such features would likely either compromise the attack’s malicious functionality, render the traffic invalid for processing, or potentially both outcomes simultaneously.
We introduce and demonstrate the effectiveness of our PS-enabled defenses, PS-guided feature selection and PS-guided feature masking, in enhancing flow-based NIDS resilience. Experimental results across various ML-based NIDS models and public datasets show that discarding or masking highly manipulatable features (high-PS features) can maintain solid detection performance while significantly reducing vulnerability to evasion adversarial attacks. Our findings confirm that PS effectively identifies flow-based NIDS features susceptible to problem-space perturbations. This novel approach leverages problem-space NIDS domain constraints as lightweight universal defense mechanisms against evasion adversarial attacks targeting flow-based ML-NIDS.
Index Terms:
Machine Learning, Evasion Adversarial Attacks, Network security, Intrusion Detection
I Introduction
@@
I-A Feature-Space vs Problem-Space Evasion Adversarial Attacks Against ML-NIDS
It is crucial to differentiate between feature-space and problem-space (real-world objects) [7] [3]. Feature-space adversarial attacks manipulate or perturb feature vectors, assuming that the attacker can directly access and alter these features; however, this assumption is often unrealistic in practical scenarios. In contrast, problem-space adversarial attacks modify or perturb actual raw input data, such as network traffic flows (e.g., adding delays) or packets (e.g., padding the payload), which attackers can typically access. Fig. 1 illustrates this distinction.
@@ I-A Feature-Space vs Problem-Space Evasion Adversarial Attacks Against ML-NIDS
(b) Two Examples of Non-Perturb-able
Features in Problem-Space Against NIDS
@@
Feature-space attacks are often impractical against NIDS due to limited attacker access to feature vectors and complications from feature correlations and network constraints [1]. In contrast, problem-space attacks are more feasible, as external attackers can modify network packets directly. These attacks typically begin with feature-space perturbations and then translate them into real-world packet modifications (Inverse Feature-Mapping [8]). As shown in Fig. 2, attackers alter network flows to perturb targeted features after feature extraction, for example, adding payload padding to change the maximum packet length or total forward packet length. However, problem-space attacks also face challenges [1, 9, 10], including maintaining malicious objective and network functionality, keeping up with model and feature extraction updates, and adhering to NIDS feature constraints.
I-B Perturb-ability of Features in Problem-Space Against NIDS
@@
N.B. Some non-perturb-able (or robust) features may be completely unmodifiable due to problem-space limitations and correlations within the NIDS domain. For example, some backward features (features describing the network flow from server to client), like the mean size of a packet in the backward direction feature, can be extremely difficult for an attacker to access. However, it is important to note that most features can be modified through problem-space manipulations. By non-perturb-able features, we specifically refer to those that cannot be perturbed in the problem-space while maintaining the attacker’s malicious aim and complying with NIDS domain constraints. For instance, changing the destination IP to manipulate its corresponding feature in the feature vector is possible. However, doing so would disrupt the flow’s network functionality and the attacker’s malicious objective, which is why we classify it as a non-perturb-able feature.
@@ I-C Motivation and Aim
Our motivation stems from the intuitive assumption that attackers can only access the problem-space rather than the feature-space. This perspective aligns with the reality of most network environments, where attackers can manipulate packet contents but do not have direct control over the feature extraction process (see section II for more details on our threat model).
@@
In response to this, our aim is to introduce the novel notion of the Perturb-ability Score (PS) metric, which is designed to enhance the robustness of ML-based NIDS. The PS metric helps to identify features in the problem-space that are susceptible to manipulation by attackers, without compromising the malicious or network functionality of traffic. By quantifying the perturb-ability of each feature within NIDS domain constraints, PS facilitates the selection of features that are inherently more resistant to adversarial attacks or the masking of nonresistant features. Our aimed classification is shown in Fig. 4. For the remainder of the paper, we will use the color scheme found in Fig. 4, i.e., a green feature represents a feature in the Low Perturb-ability class, a yellow feature represents a feature in the Medium Perturb-ability class, and a red feature represents a feature in the High Perturb-ability class.
@@
What sets our proposed defenses apart is their independence from attack types, attack norms [19] used, or the level of adversarial knowledge (whether black-box, white-box, or gray-box [3]). Unlike most conventional defenses that focus primarily on the internal mechanisms of the ML model itself, our method takes an ”outside-the-box” perspective. By leveraging inherent network domain constraints external to the ML model, we significantly reduce the attack surface. This strategic shift from model-centric defenses to exploiting network domain properties introduces a novel layer of protection, enhancing the overall robustness of ML-NIDS against adversarial threats.
I-D Contributions
@@
Our contributions are threefold: 1 Introduction of Perturb-ability Score (PS) (section III): We propose a novel Perturb-ability Score (PS) metric to quantify the vulnerability of NIDS features to adversarial manipulation. The PS measures the susceptibility of a feature to perturbation by an attacker without compromising the underlying malicious objective or network functionality of the attack. This metric establishes a robust foundation for assessing feature resilience in ML-NIDS adversarial contexts. 2 Leveraging PS in Defensive Mechanisms (section IV): By utilizing PS, we introduce defense strategies that enhance the robustness of ML-based NIDS. 2a PS-guided feature selection (section IV-A) enables the selection of only inherently resilient features that remain robust against adversarial perturbations during pre-processing. This approach fortifies NIDS by reducing the attack surface while maintaining solid detection performance across various models and datasets. 2b PS-guided feature masking (section IV-B) of high perturb-ability features, which also narrows the attack surface by replacing easily perturb-able features with neutral values, effectively removing the attacker’s ability to manipulate those features while maintaining the model’s structure and dimensionality. 3 Mapping Problem-Space Adversarial Attacks to Feature-Space (section IX) and Conducting Thorough Testing (sections VII, VIII and X): We conduct an in-depth analysis of problem-space adversarial attack techniques found in the literature by mapping their traffic morphing methods to the corresponding NIDS features. This mapping validates the PS classification, demonstrating how PS effectively captures the impact of problem-space evasion techniques on NIDS features and provides significant insights into how adversarial attacks manifest in both problem and feature-spaces. We also conducted meticulous testing and comparison using feature-space and problem-space attacks, comparisons using three public datasets, multiple models and comparisons with adversarial training.
II Threat Model
@@
Although recent research has cast doubt on the practicality of evasion adversarial attacks against NIDS [1, 21, 22], since knowing the ML algorithm, activation functions, or querying the NIDS is unlikely, this paper nonetheless focuses on exploring the possibility and implications of adding perturbations to NIDS features in the problem-space.
III Perturb-ability Score (PS)
In this section, we explain how our novel Perturbability Score (PS) quantifies the susceptibility of each flow-based NIDS feature to problem-space evasion attack perturbation.
III-A NIDS Features Perturb-ability Classification
@@ III-B PS Formulation
(a) Visualization of the PS2[$f_{i}$] Evaluation Equation, where MaxR=255 and MinR=2
@@
III-B1 PS1[$f_{i}$]: Critical Protocol, Identifiers and Functional Integrity Felids
This PS field focuses on critical features where perturbations in the problem-space could impact network functionality or enable malicious behavior in network flows.
PS1[$f_{i}$] will be 0 if any of the following conditions are true (which will make PSTotal[$f_{i}$] equals 0);
C1: the feature $f_{i}$ is protocol identification feature.
C2: the feature $f_{i}$ is a critical address/identifier (e.g., IP address in a TCP flow, port number).
C3: the feature $f_{i}$ is a field that defines a network functional integrity attribute (e.g., Service, State, flow direction)
@@ III-B1 PS1[$f_{i}$]: Critical Protocol, Identifiers and Functional Integrity Felids
III-B2 PS2[$f_{i}$]: The range of Possible Values of a Feature
@@
PS2[$f_{i}$] will be 1 if $f_{i}$’s number of Possible Values ($PV$) is greater than MaxR, where we set $\text{MaxR}=255$ (this threshold is chosen because features with this many possible values behave similarly to computer vision’s pixels, offering significant flexibility for perturbation). On the other hand, if $f_{i}$’s $PV$ ($PV$[$f_{i}$]) is less than or equal to MaxR, PS2[$f_{i}$] will be equal to a linear function where its output is 1 if $f_{i}$’s $PV$ is MaxR, and 0.5 if $f_{i}$’s $PV$ is MinR, where we set $\text{MinR}=2$ (representing binary features). If $PV[f_{i}]$ is less than MinR (equals 1), it indicates that $f_{i}$ is non-perturb-able, in which case $\text{PS}_{2}[f_{i}]$ will be set to 0. However, in this case (where $PV[f_{i}]$ equals 1), we recommend dropping that feature, as it does not contribute meaningful information to the ML model. Fig. 5a shows the visualization of the PS2[$f_{i}$] evaluation equation. It is worth noting that practitioners and domain experts can calibrate the MinR and MaxR thresholds based on specific application requirements and the nature of the feature-space under consideration.
@@
$\text{PS}_{2}[f_{i}]=\begin{cases}1&\text{if }\text{PV[$f_{i}$]}>\text{MaxR}\\ 0&\text{if }\text{PV[$f_{i}$]}<\text{MinR}\\ 0.5+(0.5\times\frac{(\text{PV[$f_{i}$]}-\text{MinR})}{(\text{MaxR}-\text{MinR})})&\text{otherwise}\end{cases}$
@@
III-B3 PS3[$f_{i}$]: Correlated Features
This PS field considers the correlation between a NIDS feature and other features. Due to network constraints within NIDS, many features exhibit problem-space correlations. For instance, the flow duration feature is typically correlated with the total forward and backward inter-arrival times. Such correlated features limit the attacker’s flexibility. The gradients of the targeted model might recommend a specific perturbation to one feature and a different perturbation to another. However, achieving these opposing perturbations simultaneously is very difficult if the features are highly correlated within the problem-space. As an example, an attacker cannot simultaneously increase the flow duration while decreasing both the forward and backward inter-arrival times. As the number of correlated features associated with a single feature increases, it becomes more difficult to perturb that feature in the problem-space. These correlations also cause collateral damage effects, which we analyze in Section IX.
When two features are correlated, there is a 50% probability that gradients will require perturbations in the same direction (either both gradients align with or oppose each other). Each additional correlated feature doubles the number of possible gradient configurations and halves the probability of the gradients requiring perturbations in a similar direction. As a result, this probability decreases exponentially as more correlated features are added. To account for this phenomenon in our PS3[$f_{i}$] calculation, we introduce an exponential formula that appropriately weights the correlation effects. Fig. 5b illustrates the PS3[$f_{i}$] evaluation equation.
PS3[$f_{i}$] can be described with the following equation, where CF[$f_{i}$] represents the number of correlated features to feature $f_{i}$:
$\text{PS}_{3}[f_{i}]=0.5+(0.5\times\frac{1}{2^{\text{CF[$f_{i}$]}}})$
We have deliberately constrained the range of values from 1 to 0.5 to ensure that even in scenarios with numerous correlated features, the PSTotal value is only slightly reduced. This design choice acknowledges that some attacks are blind in nature and do not rely on gradient information, making them less susceptible to correlation effects. The exponential formula therefore provides a balanced approach that maintains defensive integrity while accurately modeling the diminishing impact of correlation on attack success probability.
@@ III-B3 PS3[$f_{i}$]: Correlated Features
Nevertheless, we acknowledge that an attacker with exceptional knowledge of the attacked ML model and its feature vector could potentially overcome the constraints posed by correlated features in the problem-space. This is another reason why the minimal value of PS3[$f_{i}$] is set to 0.5, resulting in only a modest reduction in PStotal[$f_{i}$]. However, we still believe that correlated features in the problem-space would pose a significant challenge for an attacker in a practical setting where they have no information about the attacked ML-NIDS and cannot query it. Additionally, as the number of correlated features increases, it becomes exponentially harder for an attacker to handle these correlations.
III-B4 PS4[$f_{i}$]: Features that attackers cannot access